expansion of gases-variation with pessure

Grade:12

1 Answers

Rathod Shankar AskiitiansExpert-IITB
13 years ago

Dear vishal chintalapati,

1) for isobaric with no pressure change the volumetric thermal expansivity, which we denote βp, is:

PV = nRT \,
\ln\left(V\right) = \ln \left(T\right) + \ln\left(nR/P\right)
\beta_p = \bigg(\frac{1}{V} \frac{dV}{dT}\bigg)_p  = \bigg(\frac{d(ln V)}{d T}\bigg)_p = \frac{d(ln T)}{d T} = \frac{1}{T}



2) adibatic equation where temparature is constant
PV =Constant


Thanks
Rathod Shankar
IIT Bombay